# Stolz positive scalar curvature structure groups, proper actions and equivariant 2-types

### v2

@@ 5.2 Construction of a universal space
Let $T\colon\mathcal{G}\to\mathcal{H}$ be a morphism of groupoids.
Recall the standard two facts:
  1. 1.
@@
The assertion follows directly from the following commutative diagram
${\displaystyle\bigsqcup_{[x]\in\pi_{0}(\mathcal{G})}\mathcal{G}^{x}_{x}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}\scriptstyle{\simeq}$${\displaystyle\bigsqcup_{[x]\in\pi_{0}(\mathcal{H})}\mathcal{H}^{T(x)}_{T(x)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$${\mathcal{G}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{T}\textstyle{\mathcal{H}}$
@@
Let $X$ and $Y$ be $G$-CW-complexes and $f\colon X\to Y$ be a cellular $G$-map. Then the induced transformation
$f_{\#}\colon\Pi_{1}(X;G)\to\Pi_{1}(Y;G)$
is an equivalence in the sense of Definition 5.10 if and only if for each subgroup $H$ of $G$ and for each $x_{0}\in X^{H}$ $f$ induces isomorphisms
$\pi_{j}(f^{H})\colon\pi_{j}(X^{H},x_{0})\to\pi_{j}(Y^{H},f(x_{0}))\text{ for }j=0,1.$
@@
First, assume that $f_{\#}$ is an equivalence. By definition, this means that $f^{H}$ induces an equivalence of groupoids between $\Pi_{1}(X^{H})$ and $\Pi_{1}(Y^{H})$. This, in turn, by Remark 5.11 implies that $\pi_{j}(f^{H})$ is an isomorphism for $j=0$ and $j=1$. The other implication is a special case of Lemma 5.12. ∎
@@
Let $\Pi\colon\operatorname{Orb}(G)\to Groupoids$ be a functor (thought of as an abstract fundamental groupoid functor). Let $X$ be a $G$-CW complex and $\Phi\colon\Pi_{1}(X;G)\to\Pi$ be an equivalence as in Definition 5.10.
Then for the classifying $G$-map $\varphi\colon X\to B\Pi$ as in Proposition 5.7 which induces $\Phi$ on the level of fundamental groupoid functors, we have that for $n\geq 6$
$\varphi_{*}\colon R_{n}^{spin}(X)^{G}\rightarrow R_{n}^{spin}(B\Pi)^{G}$ (5.4)
@@
By Lemma 5.13, the map $f\colon X\to Y$ induces isomorphisms of $\pi_{0}$ and $\pi_{1}$ for each fixed point set $X^{H}$ and each basepoint. Moreover, as all higher homotopy groups of fixed point sets of $B\Pi$ are trivial, this map indeed is $2$-connected in the sense of 2.6. Therefore, the assertion is exactly the one of Theorem 4.1. ∎
Using this proposition, we arrive at a slight strengthening of Theorem 4.1:
@@
Let $f\colon X\to Y$ be a $G$-map between $G$-CW-complexes such that for each subgroup $H$ of $G$ the induced map $f^{H}\colon X^{H}\to Y^{H}$ induces an isomorphism $\pi_{j}(f^{H})\colon\pi_{j}(X^{H},x)\to\pi_{j}(Y^{H},f(x))$ for all $x\in X^{H}$ and $j=0,1$. Then for $n\geq 6$ the induced map
$f_{*}\colon R^{spin}_{n}(X)^{G}\to R^{spin}_{n}(Y)^{G}$
@@
Note that Proposition 5.14 and Corollary 5.15 of course also hold if $G$ is trivial and just state that $R_{n}^{spin}(X)$ depends only on the fundamental group of the CW-complex $X$.

### v3

@@ 5.2 Construction of a universal space
Let $T\colon\mathcal{G}\to\mathcal{H}$ be a morphism of groupoids. Recall the standard two facts:
  1. (1)
@@
The assertion follows directly from the following commutative diagram:
${\displaystyle\bigsqcup_{[x]\in\pi_{0}(\mathcal{G})}\mathcal{G}^{x}_{x}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}\scriptstyle{\simeq}$${\displaystyle\bigsqcup_{[x]\in\pi_{0}(\mathcal{H})}\mathcal{H}^{T(x)}_{T(x)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$${\mathcal{G}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{T}$${\mathcal{H}.}$
@@
Let $X$ and $Y$ be $G$-CW-complexes and $f\colon X\to Y$ be a cellular $G$-map. Then the induced transformation $f_{\#}\colon\Pi_{1}(X;G)\to\Pi_{1}(Y;G)$ is an equivalence in the sense of Definition 5.10 if and only if for each subgroup $H$ of $G$ and for each $x_{0}\in X^{H}$ $f$ induces isomorphisms
$\pi_{j}\bigl(f^{H}\bigr)\colon\ \pi_{j}(X^{H},x_{0})\to\pi_{j}\bigl(Y^{H},f(x_{0})\bigr)\qquad\text{for}\ j=0,1.$
@@
First, assume that $f_{\#}$ is an equivalence. By definition, this means that $f^{H}$ induces an equivalence of groupoids between $\Pi_{1}\bigl(X^{H}\bigr)$ and $\Pi_{1}\bigl(Y^{H}\bigr)$. This, in turn, by Remark 5.11 implies that $\pi_{j}\bigl(f^{H}\bigr)$ is an isomorphism for $j=0$ and $j=1$. The other implication is a special case of Lemma 5.12. ∎
@@
Let $\Pi\colon\operatorname{Orb}(G)\to Groupoids$ be a functor ((thought of as an abstract fundamental groupoid functor)). Let $X$ be a $G$-CW complex and $\Phi\colon\Pi_{1}(X;G)\to\Pi$ be an equivalence as in Definition 5.10. Then for the classifying $G$-map $\varphi\colon X\to B\Pi$ as in Proposition 5.7 which induces $\Phi$ on the level of fundamental groupoid functors, we have that for $n\geq 6$
$\varphi_{*}\colon R_{n}^{\rm spin}(X)^{G}\rightarrow R_{n}^{\rm spin}(B\Pi)^{G}$
@@
By Lemma 5.13, the map $f\colon X\to Y$ induces isomorphisms of $\pi_{0}$ and $\pi_{1}$ for each fixed point set $X^{H}$ and each basepoint. Moreover, as all higher homotopy groups of fixed point sets of $B\Pi$ are trivial, this map indeed is $2$-connected in the sense of Definition 2.6. Therefore, the assertion is exactly the one of Theorem 4.1. ∎
Using this proposition, we arrive at a slight strengthening of Theorem 4.1.
@@
Let $f\colon X\to Y$ be a $G$-map between $G$-CW-complexes such that for each subgroup $H$ of $G$ the induced map $f^{H}\colon X^{H}\to Y^{H}$ induces an isomorphism
$\pi_{j}\bigl(f^{H}\bigr)\colon\ \pi_{j}\bigl(X^{H},x\bigr)\to\pi_{j}\bigl(Y^{H},f(x)\bigr)$
for all $x\in X^{H}$ and $j=0,1$. Then for $n\geq 6$ the induced map
$f_{*}\colon\ R^{\rm spin}_{n}(X)^{G}\to R^{\rm spin}_{n}(Y)^{G}$
@@
Note that Proposition 5.14 and Corollary 5.15 of course also hold if $G$ is trivial and just state that $R_{n}^{\rm spin}(X)$ depends only on the fundamental group of the CW-complex $X$.